OpenStudy (anonymous):

Integral of (3x-2)/(x+1)?

OpenStudy (anonymous):

split the fraction ...

OpenStudy (anonymous):

or a u sub maybe ... let u = x+1, x = u-1 dx = du \[\int \frac{3(u-1)-2}{u}~du\] \[\int \frac{3u-5}{u}~du\]

OpenStudy (anonymous):

^.^ \[\Large \frac{3x - 2}{x+1}=\frac{3x+3}{x+1}- \frac5{x+1}=3 - \frac5{x+1}\] ;)

OpenStudy (anonymous):

yeah, that too ;)

OpenStudy (anonymous):

Thank you so much everyone